### v2

# Predicting Parkinson’s Disease Progression Using Statistical and Neural Mixed Effects Models: Comparative Study on Longitudinal Biomarkers

Abstract
Predicting Parkinson’s Disease (PD) progression is crucial, and voice biomarkers offer a non-invasive method for tracking symptom severity (UPDRS scores) through telemonitoring. Analyzing this longitudinal data is challenging due to within-subject correlations and complex, nonlinear patient-specific progression patterns. This study benchmarks LMMs against two advanced hybrid approaches: the Generalized Neural Network Mixed Model (GNMM) [17], which embeds a neural network within a GLMM structure, and the Neural Mixed Effects (NME) model [27], allowing nonlinear subject-specific parameters throughout the network. Using the Oxford Parkinson’s telemonitoring voice dataset, we evaluate these models’ performance in predicting Total UPDRS to offer practical guidance for PD research and clinical applications.
Keywords Parkinson’s Disease $\cdot$ Biostatistics $\cdot$ Longitudinal Data $\cdot$ Neural Networks $\cdot$ Artificial Intelligence
1 Introduction
Parkinson’s disease (PD) is a progressive neurodegenerative disorder marked by motor symptoms such as tremor, rigidity, and postural instability, each of which lowers quality of life. The condition affects millions worldwide, and prevalence rises with age, which makes modeling the progression of Parkinson’s disease increasingly important and urgent.
In order to measure the progression of Parkinson’s disease, researchers have explored several objective digital biomarkers. Wearable inertial sensors quantify gait impairment, bradykinesia, and tremor during everyday activity [5]. Smartphone accelerometers, gyroscopes, and touch-screen interactions capture movement patterns and tapping speed that relate to symptom severity [2]. Handwriting and drawing tasks recorded on digitizing tablets reveal micrographia and fine-motor deficits [8].
Among these methods, voice biomarkers have emerged as a promising one for tracking PD progression[23], as it is objective, non-invasive and convenient to be obtained. Subtle shifts in pitch, loudness, and stability may appear as the disease progresses, and telemonitoring makes it possible to collect frequent longitudinal voice data that complement clinic visits.These voice features are typically used to predict scores from clinical assessments like the Unified Parkinson’s Disease Rating Scale (UPDRS) [10], which serves as the primary response variable for quantifying symptom severity and progression in many PD studies.
Because the same individual is measured many times, longitudinal voice data contain within‑subject correlation, and patients differ in baseline severity and rate of change. Models therefore need to represent both population trends and subject‑specific variation [17]. Furthermore, the link between high‑dimensional voice features and UPDRS may be highly nonlinear [17].
Classical statistical methods like Linear Mixed Models (LMMs) [13] and Generalized Linear Mixed Models (GLMMs) [4] have long been common tools for analyzing such longitudinal data, as they effectively use random effects to capture within-subject correlations and individual differences. However, their primary limitation lies in the inherent assumption of linear relationships for the fixed effects component, which may inadequately model the complex, nonlinear patterns of change, which are often observed in PD progression [17].
While Nonlinear Mixed Effects (NLME) models [15] provide greater flexibility for nonlinear trends, their traditional optimization algorithms can be computationally demanding and may not scale efficiently to the high-dimensional parameter spaces characteristic of modern machine learning approaches [27]. Similarly, semi-parametric extensions like Generalized Additive Mixed Models (GAMMs) [14, 26], which use smoothing splines for time trends, can also face challenges with intricate interactions among predictors.
Deep neural networks (DNNs) offer a powerful alternative for modeling complex, nonlinear structures within large datasets. However, standard DNNs typically presuppose that observations are independent. Applying them naively to longitudinal data by disregarding these inherent correlations can result in biased estimates and suboptimal predictive performance [17]. Early adaptations, such as incorporating subject identifiers as input features in ANNs [16], aimed to address this but often encountered scalability problems as the parameter space increased with the number of subjects [17].
More recent advancements have focused on combing neural networks with mixed-effects frameworks. A prevalent strategy has been the development of Neural Networks with Linear Mixed Effects (NN-LME), where a neural network learns nonlinear data representations, and a linear mixed model is subsequently applied to these features or forms the final layer [28]. Although these NN-LME models can capture nonlinear population-level trends, they frequently restrict person-specific (random) effects to be linear and may inherit the scalability constraints of conventional LME optimization methods [27].
To overcome these limitations, models such as the Generalized Neural Network Mixed Model (GNMM) [17] were introduced. The GNMM replaces the linear fixed-effect component of a GLMM with a neural network, thereby enhancing the ability to capture nonlinear associations while retaining the GLMM structure for random effects. Further extending this method, the Neural Mixed Effects (NME) model [27] allows nonlinear person-specific parameters to be optimized at any point within the neural network architecture, offering more flexibility and scalability for modeling individual-specific nonlinear trends.
In this study, we build a benchmark by comparing these distinct strategies for predicting longitudinal Parkinson’s Disease severity from voice: (i) traditional LMMs, (ii) the Generalized Neural Network Mixed Model (GNMM) of [17], and (iii) the Neural Mixed Effects (NME) model of [27]. Using the Oxford Parkinson’s telemonitoring data from the UCI repository [20], we evaluate their ability to predict Total UPDRS and provide practical guidance for researchers and clinicians.
2 Related Work
Even though some studies have leveraged the UCI Parkinson’s Telemonitoring dataset[20] to model disease severity, many of them don’t account for the longitudinal structure in the data. For example, Eskidere et al. [9] applied various linear and nonlinear regression techniques like Support Vector Machines (SVM) and Least Squares SVM (LS-SVM) to predict UPDRS scores based on acoustic features. However, their approaches treated each observation as an independent sample, neglecting the repeated measures structure of the dataset.
Similarly, Nilashi et al. [18] proposed a hybrid system combining noise removal, clustering, and prediction methods like Adaptive Neuro-Fuzzy Inference System (ANFIS) and Support Vector Regression (SVR), but there is the absence of incorporating random effects to model individual differences in progression. Moreover, the interpretability of the model can be complicated by the combination of multiple techniques to limit the clinical utility.
Recent deep learning methods, including Convolutional Neural Networks (CNNs) based classifiers such as the work of Ananthanarayanan et al.[1], have further advanced the field. They introduced deep learning models like CNNs and Monte Carlo-Dropout CNNs to detect PD by using voice recordings. However, their models focused on classification tasks, but did not model the progression of disease over time. Also, as we know, deep learning models often act as black boxes while powerful, resulting in a challenging extraction of clinically meaningful insights.
To improve the methods, in our study, we models the longitudinal structure of the data by comparing traditional linear mixed models (LMMs) with two recent neural extensions GNMM and NME model. Relatedly, we have studied longitudinal clinical time series modeling in other settings and found that compact recurrent models can remain competitive with more complex transformer-based approaches, which supports the idea that strong temporal inductive bias can matter as much as model size [25]. By incorporating both fixed and random effects, these models are better equipped to capture both population-level trends and subject-specific variations in disease progression.
3 Methodology
3.1 Dataset
This study utilizes the UCI Parkinson’s Telemonitoring Dataset[20], which consists of longitudinal data collected from 42 patients (28 males and 14 females) diagnosed with Parkinson’s disease. These patients were in early-stage Parkinson’s disease and recruited to a six-month trial of a telemonitoring device for remote symptom progression monitoring. The recordings, comprising a range of biomedical voice measurements, were automatically collected in patients’ homes. The dataset contains a total of 5,875 records, capturing repeated measures. Each record includes 22 variables including the following information:
Category Variable Name Type Description
Demographics subject Integer Unique identifier for each subject
age Integer Age of the subject
sex Binary Subject sex (0 = male, 1 = female)
Clinical Scores total_UPDRS Continuous Total UPDRS score, linearly interpolated
motor_UPDRS Continuous Motor UPDRS score, linearly interpolated
test_time Continuous Time since recruitment (in days)
Voice Biomarkers Jitter(%) Continuous Several measures of variation in fundamental frequency
Jitter(Abs) Continuous
Jitter:RAP Continuous
Jitter:PPQ5 Continuous
Jitter:DDP Continuous
Shimmer Continuous Several measures of variation in amplitude
Shimmer(dB) Continuous
Shimmer:APQ3 Continuous
Shimmer:APQ5 Continuous
Shimmer:APQ11 Continuous
Shimmer:DDA Continuous
NHR Continuous Ratio of noise to tonal components in the voice
HNR Continuous
RPDE Continuous Nonlinear dynamical complexity measure
DFA Continuous Fractal scaling exponent
PPE Continuous Nonlinear measure of fundamental frequency variation
Table 1: Variable descriptions for the Parkinson’s Telemonitoring Dataset.
In the early stages of Parkinson’s disease (PD), many patients show noticeable changes in their speech, such as unstable pitch, uneven loudness, hoarseness, and unclear pronunciation. PD affects the brain’s ability to control the muscles which are used for speaking, including those in the throat and chest. Since telemedicine and remote health tools become more common, voice recordings have become a useful and non-invasive way to monitor how the disease changes over time. The early diagnosis can also be helped.
Many voice-based measurements are contained in the datasets and can be grouped into four types: frequency changes (Jitter), amplitude changes (Shimmer), noise features (NHR and HNR), and nonlinear patterns (RPDE, DFA, and PPE).
Jitter features including Jitter (%), Jitter (Abs), RAP, PPQ5, and DDP measure small changes in pitch between voice cycles. Since the vocal folds of people with PD do not move smoothly, they cannot keep a steady pitch which will lead to a higher jitter values.
Shimmer feature such as Shimmer in percent and decibels, APQ3, APQ5, APQ11, and DDA show how the loudness of the voice changes from one cycle to the next. Due to the muscle control problems which make their speech volume less steady, these values are resulted in higher.
NHR (Noise-to-Harmonics Ratio) and HNR (Harmonics-to-Noise Ratio) are used to check how much noise is in the voice. PD patients tend to have more noise and less clear voice sounds which make these values worse than healthy individuals.
Furthermore, Recurrence Period Density Entropy (RPDE), Detrended Fluctuation Analysis (DFA), and Pitch Period Entropy (PPE) are nonlinear dynamic measures that capture complexity and unpredictability in vocal patterns. People with PD may speak in a way which is less regular or harder to predict. These features can be helpful to identify those subtle inssues in voice control.
Overall, these voice features can help to study speech problems in Parkinson’s disease. These features can be easily measured and also related to related to motor symptoms closely, therefore they are valuable digital biomarkers in tracking disease progression and supporting remote healthcare systems.
In this study, we use total_UPDRS as the response variable to explore the impact of voice biomarkers on disease progression.
(a) Correlation heatmap
(b) Scatter plot of total_UPDRS versus age
(c) Subjectspecific log(total_UPDRS) trajectories over time
Figure 1: Data analysis of the dataset
Figure 1(a) shows the pairwise Pearson correlations among all measured variables. Several voice-based features, particularly jitter and shimmer measures, show moderate positive correlations with UPDRS scores. This indicates that voice instability is related to the severity of the disease.
Figure 1(b) illustrates the relationship between patient age and total_UPDRS scores. Although the scatterplot reveals variability within the age range, the fitted regression line shows a clear positive association, indicating that older individuals tend to have more severe symptoms.
Figure 1(c) presents longitudinal trends in log-transformed total_UPDRS scores for individual patients. The patterns differ among patients, with some showing progressive worsening while others remain stable or even slightly improve. The heterogeneity in progression patterns shows the necessity for individualized modeling approaches like mixed effects models.
3.2 Traditional Methods: Linear Mixed–Effects Model (LMM)
We first apply Linear Mixed-Effects model [13] on our dataset. Let $Y_{ij}$ denote the UPDRS score for subject $i\;(i=1,\dots,m)$ at time $t_{ij}\;(j=1,\dots,n_{i})$, and let $X_{ijk}$ be the kk‑th voice feature ($k=1,\dots,K$). We use a linear mixed–effects model with a subject‑specific random intercept $b_{0i}$ and random slope $b_{1i}$ for time:
$Y_{ij}\;=\;\beta_{0}\;+\;\beta_{1}t_{ij}\;+\;\sum_{k=1}^{K}\beta_{k+1}\,X_{ijk}\;+\;b_{0i}\;+\;b_{1i}t_{ij}\;+\;\varepsilon_{ij}.$ (1)
In this model, the terms $\beta_{0}$, $\beta_{1}t_{ij}$, and $\sum_{k=1}^{K}\beta_{k+1}\,X_{ijk}$ represent the fixed effects. Specifically, $\beta_{0}$ is the overall intercept, $\beta_{1}$ is the average slope for time $t_{ij}$ across all subjects, and $\beta_{k+1}$ are the coefficients for the $K$ voice features $X_{ijk}$, representing their average effects on $Y_{ij}$. These fixed effects describe the population-average relationships.
The terms $b_{0i}$ and $b_{1i}t_{ij}$ represent the random effects for subject $i$. Here, $b_{0i}$ is the subject-specific random intercept, showing how subject $i$’s baseline UPDRS score deviates from the overall intercept $\beta_{0}$. Similarly, $b_{1i}$ is the subject-specific random slope for time, indicating how subject $i$’s rate of change in UPDRS score over time $t_{ij}$ deviates from the average time slope $\beta_{1}$. These random effects capture individual heterogeneity around the population-average trends.
Finally, $\varepsilon_{ij}$ is the residual error term for subject $i$ at time $j$, representing within-subject variability not explained by the fixed or random effects.
Distributional assumptions.
The subject-specific random effects $(b_{0i},b_{1i})^{\top}$ are assumed to be drawn from a bivariate normal distribution with a mean of zero and a covariance matrix $D$:
$\begin{pmatrix}b_{0i}\\ b_{1i}\end{pmatrix}\,\sim\,\mathcal{N}\!\Bigl(\mathbf{0},\,D=\begin{pmatrix}\sigma_{b0}^{2}&\rho\,\sigma_{b0}\sigma_{b1}\\ \rho\,\sigma_{b0}\sigma_{b1}&\sigma_{b1}^{2}\end{pmatrix}\Bigr)$.
The residual errors $\varepsilon_{ij}$ are assumed to be independent and identically distributed (i.i.d.) normal random variables with a mean of zero and variance $\sigma^{2}$:
$\varepsilon_{ij}\sim\mathcal{N}(0,\sigma^{2}).$
Furthermore, the random effects $\mathbf{b}_{i}$ and residual errors $\bm{\varepsilon}_{i}$ are assumed to be independent of each other.
Matrix formulation.
Let $\mathbf{y}_{i}=(Y_{i1},\dots,Y_{in_{i}})^{\top}$ be the vector of $n_{i}$ UPDRS scores for subject $i$. The fixed effects design matrix $\mathbf{X}_{i}$ and the random effects design matrix $\mathbf{Z}_{i}$ for subject $i$ are defined as:
$\mathbf{X}_{i}=\begin{pmatrix}1&t_{i1}&X_{i1,1}&\dots&X_{i1,K}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&t_{in_{i}}&X_{in_{i},1}&\dots&X_{in_{i},K}\end{pmatrix},\quad\mathbf{Z}_{i}=\begin{pmatrix}1&t_{i1}\\ \vdots&\vdots\\ 1&t_{in_{i}}\end{pmatrix}$.
The matrix $\mathbf{X}_{i}$ contains a column of ones for the intercept, a column for time $t_{ij}$, and $K$ columns for the voice features $X_{ijk}$. The matrix $\mathbf{Z}_{i}$ contains a column of ones for the random intercept and a column for time $t_{ij}$ corresponding to the random slope. Then the model for subject $i$ can be written as:
$\mathbf{y}_{i}\;=\;\mathbf{X}_{i}\bm{\beta}\;+\;\mathbf{Z}_{i}\mathbf{b}_{i}\;+\;\bm{\varepsilon}_{i},\qquad\mathbf{b}_{i}\sim\mathcal{N}(\mathbf{0},D),\;\bm{\varepsilon}_{i}\sim\mathcal{N}(\mathbf{0},\sigma^{2}I_{n_{i}}).$
where $\bm{\beta}=(\beta_{0},\beta_{1},\dots,\beta_{K+1})^{\top}$ is the vector of fixed-effects coefficients, $\mathbf{b}_{i}=(b_{0i},b_{1i})^{\top}$ is the vector of random effects for subject $i$, and $\bm{\varepsilon}_{i}=(\varepsilon_{i1},\dots,\varepsilon_{in_{i}})^{\top}$ is the vector of residual errors for subject $i$.
The marginal distribution is
$\mathbf{y}_{i}\;\sim\;\mathcal{N}\!\bigl(\mathbf{X}_{i}\bm{\beta},\;\mathbf{V}_{i}\bigr),\qquad\mathbf{V}_{i}\;=\;\mathbf{Z}_{i}D\mathbf{Z}_{i}^{\top}+\sigma^{2}I_{n_{i}}.$
Stacking all subjects gives $\mathbf{y}\sim\mathcal{N}(\mathbf{X}\bm{\beta},\mathbf{V})$ with $\mathbf{V}=\mathrm{blockdiag}(\mathbf{V}_{1},\dots,\mathbf{V}_{m})$.
The log‑likelihood can be written as:
$\ell(\bm{\beta},\theta)\;=\;-\tfrac{1}{2}\Bigl\{\log|\mathbf{V}|+(\mathbf{y}-\mathbf{X}\bm{\beta})^{\top}\mathbf{V}^{-1}(\mathbf{y}-\mathbf{X}\bm{\beta})+n\log(2\pi)\Bigr\},$ (2)
where $\theta=(\sigma_{b0}^{2},\sigma_{b1}^{2},\rho,\sigma^{2})$ represents the vector of variance components. The estimation of $\bm{\beta}$ and $\theta$ via Maximum Likelihood (ML) proceeds as follows. Setting $\partial\ell/\partial\bm{\beta}=\mathbf{0}$ gives the generalised least‑squares (GLS) estimator for $\bm{\beta}$, conditional on $\theta$:
$\hat{\bm{\beta}}\;=\;\bigl(\mathbf{X}^{\top}\mathbf{V}^{-1}\mathbf{X}\bigr)^{-1}\mathbf{X}^{\top}\mathbf{V}^{-1}\mathbf{y}.$ (3)
For each variance component $\theta_{j}$ in $\theta$, the ML estimate $\hat{\theta}_{j}$ is found by solving the score equation $\partial\ell/\partial\theta_{j}=\mathbf{0}$:
$\frac{\partial\ell}{\partial\theta_{j}}\;=\;-\tfrac{1}{2}\Bigl\{\operatorname{tr}\!\bigl(\mathbf{V}^{-1}\partial_{\theta_{j}}\mathbf{V}\bigr)-(\mathbf{y}-\mathbf{X}\bm{\beta})^{\top}\mathbf{V}^{-1}(\partial_{\theta_{j}}\mathbf{V})\mathbf{V}^{-1}(\mathbf{y}-\mathbf{X}\bm{\beta})\Bigr\}=0.$
These equations are typically solved numerically (e.g., Newton–Raphson), often by iterating between estimating $\bm{\beta}$ given $\theta$, and then $\theta$ given $\bm{\beta}$, until convergence to obtain $\hat{\bm{\beta}}$ and $\hat{\theta}$.
While ML provides estimates for all parameters, its estimates of variance components ($\hat{\theta}$) can be biased, particularly in smaller samples, as ML does not fully account for the degrees of freedom used to estimate the fixed effects ($\bm{\beta}$).
Restricted Maximum Likelihood (REML) is preferred for estimating variance components as it yields less biased estimates. REML achieves this by maximizing a likelihood function based on linear combinations of $\mathbf{y}$ that are invariant to the fixed effects, effectively adjusting for the estimation of $\bm{\beta}$. The REML log-likelihood is:
$\ell_{\text{REML}}(\theta)\;=\;-\tfrac{1}{2}\Bigl\{\log|\mathbf{V}|+\log|\mathbf{X}^{\top}\mathbf{V}^{-1}\mathbf{X}|+(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})^{\top}\mathbf{V}^{-1}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})+(n-p)\log(2\pi)\Bigr\},$ (4)
where $p=\dim(\bm{\beta})$, and $\hat{\bm{\beta}}$ is the GLS estimator from (3). REML estimates $\hat{\theta}_{\text{REML}}$ are found by solving $\partial\ell_{\text{REML}}/\partial\theta_{j}=\mathbf{0}$ numerically. Subsequently, $\bm{\beta}$ is estimated using GLS with $\mathbf{V}$ evaluated at $\hat{\theta}_{\text{REML}}$.
Numerical Estimation in Practice
Closed‑form solutions for the variance components $\theta$ do not exist, so one resorts to iterative algorithms. Two standard choices are (i) Newton / Fisher scoring on the log‑likelihood and (ii) the expectation–maximisation (EM) algorithm that treats the random effects $\mathbf{b}$ as latent variables, which can be implemented with standard software (e.g. lme4 in R or statsmodels in Python).
Variable Selection
In modeling Parkinson’s disease progression, the presence of multicollinearity among voice biomarkers poses a significant challenge. Several acoustic features, such as the various jitter and shimmer measures, are known to be highly correlated (e.g., as shown in Figure 1 (a), correlations exceeding 0.9 between some shimmer metrics). To address this and ensure model interpretability and parsimony, we employed a two-stage variable selection strategy: first using the Least Absolute Shrinkage and Selection Operator (LASSO) for dimensionality reduction, followed by backward stepwise selection on a linear mixed-effects model. Variance Inflation Factor (VIF) diagnostics were also computed to assess residual multicollinearity after selection.
The LASSO was applied on the linear model ignoring random effects, focusing solely on the fixed effects which shrinks less informative coefficients to zero, yielding a sparse set of candidate predictors. Importantly, the LASSO helped identify redundant jitter and shimmer variables, retaining only the most informative features for further modeling.
Subsequently, we performed stepwise backward selection using the lmer model, starting from a full linear mixed-effects model with all LASSO-selected predictors. This iterative process removed non-significant fixed effects based on AIC, leading to a reduced yet effective model. During this step, we also checked VIF values to confirm that no remaining variable exhibited severe multicollinearity (all VIF <5.0<5.0).
The final model retained six predictors: age, test_time, Jitter_PPQ5, NHR, and HNR. This subset balances interpretability, predictive performance, and model stability, and serves as the foundation for further modeling, including transformation, interaction terms, and random slopes. Note that interaction effects and nonlinear terms were considered in later modeling stages rather than during variable selection.
Model Refinement via Interaction and Random Slopes
After identifying a subset of relevant predictors through variable selection, we further refined the linear mixed-effects model by incorporating interaction terms and evaluating the inclusion of subject-specific random slopes. These enhancements were designed to capture individual variation more flexibly and to model potential time-varying effects, thereby improving overall model fit and predictive performance.
To ensure the validity of the model assumptions, we examined diagnostic plots of residuals and normality. Figure 2 (top row) presents diagnostic plots from the initial model using the raw outcome variable total_UPDRS. The residual plot reveals heteroscedasticity, with increasing spread at higher fitted values, while the Q-Q plots for both fixed and random effects show noticeable deviations from normality.
To mitigate these issues, we applied a logarithmic transformation to the outcome variable. This transformation significantly stabilized the variance and improved the normality of residuals, as shown in the bottom row of Figure 2. It also compressed the scale of extreme values, reducing the influence of high-leverage tail points and producing more symmetric residuals overall.
Figure 2: Top row: Residuals, fixed effect Q-Q, and random effect Q-Q plots from the original model using total_UPDRS. Bottom row: Diagnostics after log-transforming the response. Transformation improves variance stabilization and normality.
In addition to transforming the response, we explored potential interactions among the selected covariates. A systematic evaluation of all pairwise interactions using likelihood-based model comparison revealed that the interaction between test_time and HNR was the most impactful. This interaction was statistically significant and led to a notable improvement in model fit, with the AIC decreasing from 10231.6-10231.6 to 10243.3-10243.3. The result suggests that the effect of HNR on disease progression varies over time. Other interactions provided marginal improvement or introduced unnecessary complexity and were therefore excluded from the final model.
We also investigated whether to include random slopes in addition to random intercepts for each subject. As shown in Figure 1 (c), subject-specific log(UPDRS) trajectories over time exhibited heterogeneous slopes, motivating the inclusion of a random slope for test_time. The addition of this random slope further reduced AIC to 10261.4-10261.4, resulting in the final model:
$\log(\text{UPDRS}*{ij})$ $=\beta_{0}+\beta_{1}\text{age}_{i}+\beta_{2}\text{test\_time}_{ij}+\beta_{3}\text{HNR}_{ij}+\beta_{4}(\text{test\_time}_{ij}\times\text{HNR}_{ij})$
$+b_{0i}+b_{1i}\text{test\_time}_{ij}+\varepsilon_{ij}$
where $b_{0i},b_{1i}\sim\mathcal{N}(0,G)$ represent the subject-specific random intercept and slope for time, and $\varepsilon_{ij}\sim\mathcal{N}(0,\sigma^{2})$ denotes the residual error.
Table 2 summarizes the model refinement process using AIC as the selection criterion. Each modification, i.e., the selection of variables, the transformation, the inclusion of interactions and the random slope, improved the model fit, culminating in the final model specification above.
Table 2: Model selection and refinement steps based on AIC comparison.
Step AIC Comment
Full model (all predictors) 28124.1 Initial LMM
After LASSO (fixed effects only) 27880.6 Removed highly correlated terms
After stepwise (LMM) & VIF 27869.8 Dropped non-significant effects
After log-transformation –10231.6 Improved residual normality, variance
Add interaction (test_time $\times$ HNR) –10243.3 Included significant time-varying HNR effect
Add random slope –10261.4 Final model with varying subject-specific slopes
This finalized model serves as a baseline for comparison against more flexible methods, such as generalized additive and deep mixed-effects models, in subsequent sections.
3.3 Generalized Additive Mixed Model (GAMM)
To capture nonlinear temporal effects in Parkinson’s disease progression, we adopted a Generalized Additive Mixed Model (GAMM), which extends the linear mixed-effects framework by allowing smooth, data-driven functions of continuous covariates [26]. This formulation maintains the interpretability of linear effects while introducing the flexibility necessary to model nonlinear trends over time.
In our application, the log-transformed total UPDRS score is modeled as a smooth function of test_time, along with linear terms for other covariates. The model is expressed as:
$\log(\text{UPDRS}_{ij})$ $=\beta_{0}+\beta_{1}\text{age}_{i}+\beta_{2}\text{HNR}_{ij}+f(\texttt{test\_time}_{ij})+b_{0i}+b_{1i}\texttt{test\_time}_{ij}+\varepsilon_{ij},$ (5)
where $f(\cdot)$ is a smooth function of time, and $b_{0i},b_{1i}\sim\mathcal{N}(0,G)$ are the subject-specific random intercept and slope. The residual errors $\varepsilon_{ij}\sim\mathcal{N}(0,\sigma^{2})$ are assumed to be independent and homoscedastic.
Spline Basis Representation and Estimation
The smooth function $f(\texttt{test\_time})$ is approximated via a linear combination of basis functions:
$f(\texttt{test\_time})=\sum_{k=1}^{K}\alpha_{k}B_{k}(\texttt{test\_time}),$ (6)
where $B_{k}(\cdot)$ are predefined spline basis functions (e.g., cubic regression splines, B-splines, or thin plate splines [21]), and $\alpha_{k}$ are coefficients estimated from the data. To control smoothness and avoid overfitting, a roughness penalty is imposed on the second derivative of the function:
$\text{Penalized log-likelihood}=\ell(\bm{\beta},\bm{\alpha})-\frac{1}{2}\lambda\int\left[f^{\prime\prime}(t)\right]^{2}dt,$ (7)
where $\lambda$ is a smoothing parameter that balances model fit and regularity. The complexity of the spline is quantified via its effective degrees of freedom (edf).
Estimation proceeds using Penalized Iteratively Reweighted Least Squares (P-IRLS), an efficient approach for maximizing the penalized likelihood. The gamm() function in the mgcv R package is used to jointly estimate the fixed effects, the smooth term, and the random effects. Internally, gamm() delegates the random effects estimation to the lme() function from the nlme package, allowing for flexible modeling of subject-specific deviations via random intercepts and slopes.
Smoothing parameters $\lambda$ are selected by optimizing the marginal Restricted Maximum Likelihood (REML) criterion, which balances model fit with smoothness and has been shown to offer stability and efficiency in practice [24].
Comparison of LMM and GAMM
To assess model performance, we compared the final Linear Mixed-Effects Model (LMM) and the Generalized Additive Mixed Model (GAMM) using both model fit criteria and predictive accuracy. Table 3 summarizes the estimated fixed and random effects, along with model fit and test set performance.
The GAMM demonstrated a superior model fit based on Akaike Information Criterion (AIC), achieving a lower AIC value (16162.05-16162.05) than the LMM (16062.39-16062.39). This improvement can be attributed to GAMM’s flexibility in capturing nonlinear structures, particularly through a spline term applied to test_time. As shown in Figure 3, the estimated smooth function of time deviates notably from linearity, reinforcing the presence of nonlinear progression patterns in UPDRS scores over time.
To compare predictive performance, we conducted a hold-out evaluation: the last observation from each of the 42 subjects was set aside as the test set, while the remaining data were used for training. On this test set, GAMM achieved a slightly lower mean squared error (MSE = 6.56) compared to LMM (MSE = 7.70), indicating marginally better prediction accuracy in the original scale.
Notably, the interaction term test_time $\times$ HNR was statistically significant in the LMM but became non-significant under GAMM. This suggests that the nonlinear main effect of test_time in GAMM may account for variation previously explained by the interaction term in the LMM. The estimated standard deviations of the random effects and residual terms were similar across models, indicating consistent subject-level variation.
Table 3: Comparison of LMM and GAMM Estimates and Performance
Fixed Effects (Est. (p-val)) LMM GAMM Notes
Intercept 3.316 (<<2e-16) 3.316 (<<2e-16)
Age 0.136 (0.0219) 0.136 (0.0145) Significant in both
test_time 0.0358 (0.0099) spline Nonlinear in GAMM
HNR 0.0036 (0.0035) 0.0043 (0.0006) Consistent positive effect
test_time:HNR -0.0077 (1e-08) n.s. Only significant in LMM
Smooth Terms (GAMM only) s(test_time): edf = 6.17, p << 2e-16
Random Effects (Std. Dev (Corr))
Intercept 0.382 0.373 Similar
Slope (test_time) 0.085 (0.08-0.08) 0.084 (0.061-0.061) Similar
Residual 0.058 0.057 Similar
Model Fit (AIC) -16062.39 -16162.05 GAMM better (fit)
Test MSE (orig. scale) 7.70 6.56 GAMM better (predictive)
Figure 3: Estimated spline effect of test_time from GAMM, showing nonlinear progression over time.
In summary, GAMM provided a more flexible fit by capturing nonlinear temporal patterns, as evident in its superior AIC and spline visualization. However, LMM offered comparable generalization performance on the test set and more interpretable fixed effects. The choice between models thus depends on whether the goal prioritizes interpretability or modeling flexibility.
3.4 Generalized Neural Network Mixed Model (GNMM) for Non‑linear Longitudinal Modeling
Building on Mandel [17] we use a Generalized Neural Network Mixed Model (GNMM) to predict the longitudinal Total_UPDRS scores collected in the tele‑monitoring study of Parkinson’s disease.
We retain the notation introduced earlier: $Y_{ij}$ is the Total_UPDRS score for subject $i$ at visit $j$, $\mathbf{X}_{ij}\in\mathbb{R}^{p}$ is the p=17p=17‑vector of predictors (test time + 16 voice features).
Let $i=1,\dots,m$ label the $m=42$ patients in the Oxford telemonitoring study and $j=1,\dots,n_{i}$ their successive visits. At visit $j$ we record the response $Y_{ij}$ (the Total_UPDRS score) and a predictor vector
$\mathbf{X}_{ij}=(\texttt{test\_time},\text{16 voice features})^{\!\top}\in\mathbb{R}^{17},$
where test_time is the elapsed study time and the remaining 16 entries are acoustic measures extracted from the voice recording.
Following the mixed‑effects formulation of [17] , we allow observations from the same subject to be correlated through a cluster‑specific random‑effect vector $\mathbf{b}_{i}\in\mathbb{R}^{q}$, where $q\geq 1$. Conditional on $\mathbf{b}_{i}$, the outcomes $Y_{ij}$ are assumed to follow an exponential‑family distribution
$E\!\bigl[Y_{ij}\mid\mathbf{b}_{i}\bigr]=\mu_{ij}^{\mathbf{b}},\qquad\operatorname{Var}\!\bigl(Y_{ij}\mid\mathbf{b}_{i}\bigr)=\phi\,a_{ij}\,v\!\bigl(\mu_{ij}^{\mathbf{b}}\bigr),$
with known variance function $v(\cdot)$, fixed dispersion $\phi$, and $a_{ij}$ is a known constant.
Generalized Neural Network Mixed Model (GNMM) On Our Case
Consider a feed‑forward artificial neural network (ANN) with $L$ hidden layers, the predictor vector $\mathbf{X}_{ij}\in\mathbb{R}^{p}$ ($p=17$) as input, and a univariate output $\mu^{\mathbf{b}}_{ij}$ representing the conditional mean of $Y_{ij}$ (Total_UPDRS) for subject $i$ at visit $j$. Following [17], the network output is built up through a sequence of nested activation functions $g_{\ell}(\cdot)$, $\ell=0,\dots,L$.
Network layers.
The input $\mathbf{X}_{ij}$ enters the LL‑th (bottom) hidden layer with $k_{L}$ nodes:
$\bm{\alpha}^{(L)}_{ij}=g_{L}\!\bigl\{\bm{\omega}^{(L)}\mathbf{X}_{ij}+\bm{\delta}^{(L)}\bigr\},$ (8)
where $\bm{\omega}^{(L)}$ is a $k_{L}\times p$ weight matrix and $\bm{\delta}^{(L)}$ is a bias vector of length $k_{L}$. For hidden layer $\ell=L-1,\dots,1$ with $k_{\ell}$ nodes,
$\bm{\alpha}^{(\ell)}_{ij}=g_{\ell}\!\bigl\{\bm{\omega}^{(\ell)}\bm{\alpha}^{(\ell+1)}_{ij}+\bm{\delta}^{(\ell)}\bigr\},$ (9)
with $\bm{\omega}^{(\ell)}$ of size $k_{\ell}\times k_{\ell+1}$ and $\bm{\delta}^{(\ell)}\in\mathbb{R}^{k_{\ell}}$.
Output layer and random effects.
The univariate network output determines the conditional mean through
$\mu^{\mathbf{b}}_{ij}=g_{0}\!\bigl\{\bm{\omega}^{(0)}\bm{\alpha}^{(1)}_{ij}+\delta^{(0)}+\mathbf{Z}_{ij}^{\top}\mathbf{b}_{i}\bigr\},$ (10)
where $\bm{\omega}^{(0)}$ is a $1\times k_{1}$ weight vector, $\delta^{(0)}$ a scalar bias, $\mathbf{Z}_{ij}\in\mathbb{R}^{q}$ the design vector for the cluster‑specific random effect, and $\mathbf{b}_{i}\sim N(\mathbf{0},\mathbf{D})$. In a classical generalized linear mixed model [4],
$E\!\bigl[Y_{ij}\mid\mathbf{b}_{i}\bigr]=h\!\bigl(\mathbf{X}_{ij}^{\top}\bm{\alpha}+\mathbf{Z}_{ij}^{\top}\mathbf{b}_{i}\bigr),$ (11)
where $h(\cdot)$ is the inverse link, $\mathbf{X}_{ij}^{\top}\bm{\alpha}$ the fixed‑effect component, and $\mathbf{Z}_{ij}^{\top}\mathbf{b}_{i}$ the random effect. In the GNMM we replace the fixed‑effect term by the nonlinear network output and use the final activation $g_{0}(\cdot)$ in place of $h(\cdot)$:
$E\!\bigl[Y_{ij}\mid\mathbf{b}_{i}\bigr]=g_{0}\!\bigl(\bm{\omega}^{(0)}\bm{\alpha}^{(1)}_{ij}+\delta^{(0)}+\mathbf{Z}_{ij}^{\top}\mathbf{b}_{i}\bigr).$ (12)
Quasi‑likelihood.
Let $\bm{\omega}=\operatorname{vec}\!\bigl(\bm{\omega}^{(0)},\bm{\omega}^{(1)}\bigr)$ and $\bm{\delta}=\bigl(\delta^{(0)},\bm{\delta}^{(1)}\bigr)$ collect all weights and biases of the single‑hidden–layer network. With $q=1$ (a random intercept) we set $\mathbf{Z}_{ij}\equiv 1$ and $b_{i}\sim N(0,D)$. The quasi‑likelihood used to estimate $\bigl(\bm{\omega},\bm{\delta},D\bigr)$ is
$\exp\!\{ql(\bm{\omega},\bm{\delta},\theta)\}\;\propto\;|D|^{-1/2}\int\exp\Biggl\{\frac{1}{\phi}\sum_{i=1}^{42}\sum_{j=1}^{n_{i}}\int_{Y_{ij}}^{\mu^{b}_{ij}}\frac{y_{ij}-u}{a_{ij}v(u)}\,\mathrm{d}u\;-\;\frac{b^{2}}{2D}\;-\;\lambda\!\bigl(\bm{\omega}^{\!\top}\bm{\omega}+\bm{\delta}^{\!\top}\bm{\delta}\bigr)\Biggr\}\,\mathrm{d}b.$ (13)
Laplace approximation.
Following [4] and [17], write
$\kappa(b)=-\sum_{j=1}^{n_{i}}\!\int_{Y_{ij}}^{\mu^{b}_{ij}}\frac{y_{ij}-u}{\phi\,a_{ij}v(u)}\,\mathrm{d}u+\frac{b^{2}}{2D}+\frac{\lambda}{2}\bigl(\bm{\omega}^{\!\top}\bm{\omega}+\bm{\delta}^{\!\top}\bm{\delta}\bigr),$
so that (13) is proportional to $|D|^{-1/2}\!\int\!\exp\!\{-\kappa(b)\}\,\mathrm{d}b$. Let $\tilde{b}$ be the mode obtained from $\partial\kappa/\partial b=0$; first‑ and second‑order derivatives give
$\kappa^{\prime}(b)=-\sum_{j=1}^{n_{i}}\frac{(Y_{ij}-\mu^{b}_{ij})g_{0}^{\prime}\!\bigl(\eta_{ij}^{b}\bigr)}{\phi\,a_{ij}v\!\bigl(\mu^{b}_{ij}\bigr)}+\frac{b}{D},\qquad\kappa^{\prime\prime}(b)=Z_{i}^{\top}W_{i}Z_{i}+D^{-1},$
where $Z_{i}=\mathbf{1}_{n_{i}}$ and $W_{i}=\operatorname{diag}\!\bigl\{\phi^{-1}a_{ij}^{-1}v\!\bigl(\mu^{\tilde{b}}_{ij}\bigr)^{-1}g_{0}^{\prime}\!\bigl(\eta_{ij}^{\tilde{b}}\bigr)^{2}\bigr\}$. Ignoring the remainder term yields the Laplace approximation of the Quasi-likelihood function, which will be used in the following steps.
$ql(\bm{\omega},\bm{\delta},\theta)\;\approx\;-\frac{1}{2}\log|D|-\frac{1}{2}\log\bigl(Z_{i}^{\top}W_{i}Z_{i}+D^{-1}\bigr)-\frac{\lambda}{2}\bigl(\bm{\omega}^{\!\top}\bm{\omega}+\bm{\delta}^{\!\top}\bm{\delta}\bigr).$ (14)
Training objective.
Treating the $W_{i}$ term in (14) as negligible (as in [17]) leads to the objective maximised during network training:
$ql(\bm{\omega},\bm{\delta},\theta)\;\propto\;\frac{1}{\phi}\sum_{i=1}^{42}\sum_{j=1}^{n_{i}}\int_{Y_{ij}}^{\mu^{\tilde{b}}_{ij}}\frac{y_{ij}-u}{a_{ij}v(u)}\,\mathrm{d}u\;-\;\frac{\tilde{b}^{2}}{2D}\;-\;\lambda\!\bigl(\bm{\omega}^{\!\top}\bm{\omega}+\bm{\delta}^{\!\top}\bm{\delta}\bigr).$ (15)
Quasi‑score equations (single hidden layer).
Differentiating (15) with respect to the network parameters yields the quasi‑score system
$\frac{\partial ql}{\partial\omega^{(0)}_{k}}$ $=\frac{1}{\phi}\sum_{i=1}^{42}\sum_{j=1}^{n_{i}}(Y_{ij}-\mu^{\tilde{b}}_{ij})\,g_{0}^{\prime}\!\bigl(\eta_{ij}^{\tilde{b}}\bigr)\,\alpha^{(1)}_{ij,k}-2\lambda\,\omega^{(0)}_{k}$, (QS1)
$\frac{\partial ql}{\partial\delta^{(0)}}$ $=\frac{1}{\phi}\sum_{i=1}^{42}\sum_{j=1}^{n_{i}}(Y_{ij}-\mu^{\tilde{b}}_{ij})\,g_{0}^{\prime}\!\bigl(\eta_{ij}^{\tilde{b}}\bigr)-2\lambda\,\delta^{(0)},$ (QS2)
$\frac{\partial ql}{\partial\omega^{(1)}_{lk}}$ $=\frac{1}{\phi}\sum_{i=1}^{42}\sum_{j=1}^{n_{i}}(Y_{ij}-\mu^{\tilde{b}}_{ij})\,g_{0}^{\prime}\!\bigl(\eta_{ij}^{\tilde{b}}\bigr)\,\omega^{(0)}_{k}\,g_{1}^{\prime}\!\bigl(s_{ij,k}\bigr)\,X_{ij,l}-2\lambda\,\omega^{(1)}_{lk}$, (QS3)
$\frac{\partial ql}{\partial\delta^{(1)}_{k}}$ $=\frac{1}{\phi}\sum_{i=1}^{42}\sum_{j=1}^{n_{i}}(Y_{ij}-\mu^{\tilde{b}}_{ij})\,g_{0}^{\prime}\!\bigl(\eta_{ij}^{\tilde{b}}\bigr)\,\omega^{(0)}_{k}\,g_{1}^{\prime}\!\bigl(s_{ij,k}\bigr)-2\lambda\,\delta^{(1)}_{k}$. (QS4)
Here $\eta_{ij}^{\tilde{b}}=\bm{\omega}^{(0)}\bm{\alpha}^{(1)}_{ij}+\delta^{(0)}+\tilde{b}_{i}$, $s_{ij,k}=(\bm{\omega}^{(1)}_{k})^{\!\top}\mathbf{X}_{ij}+\delta^{(1)}_{k}$, and $\alpha^{(1)}_{ij,k}=g_{1}(s_{ij,k})$. Equation (15) is maximised by solving (QS1)–(QS4) jointly with $\kappa^{\prime}(\tilde{b})=0$; we use back‑propagation for the network parameters and stochastic gradient descent on $\tilde{b}$, updating $\tilde{b}$ at each epoch while keeping $D$ fixed at its REML estimate (as in GLMM, with $\mathbf{X}_{ij}^{\top}\bm{\beta}$ replaced by $\bm{\omega}^{(0)}\bm{\alpha}^{(1)}_{ij}+\delta^{(0)}$).
Summary of the Algorithm
The GNMM network transforms the 17‑dimensional vector of voice features into a latent disease score, allowing nonlinear interactions and saturation effects to influence the predicted Total_UPDRS. The random intercept $b_{i}$ absorbs persistent patient‑level deviations, so estimates borrow strength across subjects (partial pooling). An $L_{2}$ penalty $\lambda(\bm{\omega}^{\top}\bm{\omega}+\bm{\delta}^{\top}\bm{\delta})$ decreases large weights and biases, mitigating over‑fitting.
The algorithm flow used is summarzied as below:
Algorithm 1 Stochastic training of GNMM on the Parkinson data (adapted from [17])
1:scaled features $\mathbf{X}^{\prime}$, scaled targets $Y^{\prime}$, subject indices; epochs $E$, batch size $B$, learning rate $\eta$
2:Initialise parameters $\bm{\vartheta}$ (Xavier), $b_{i}\leftarrow 0$, $\sigma^{2}\leftarrow 1$, $\sigma_{b}^{2}\leftarrow 1$
3:for $e=1,\dots,E$ do
4:  Shuffle the training set
5:  for each mini‑batch $\mathcal{B}$ of size $B$ do
6:   Compute $\mu_{ij}^{\,b}$ for $(i,j)\in\mathcal{B}$ via (10)
7:   Evaluate mini‑batch loss $L_{\mathcal{B}}$ from ((15)
8:   Back‑propagate $\nabla_{\bm{\vartheta}}L_{\mathcal{B}}$ and $\nabla_{b}L_{\mathcal{B}}$
9:   Update $\bm{\vartheta}\leftarrow\bm{\vartheta}-\eta\,\nabla_{\bm{\vartheta}}L_{\mathcal{B}}$
10:   Update each $b_{i}$ that appears in $\mathcal{B}$
11:  end for
12:  $\sigma^{2}\leftarrow$ mean squared residual over the full training set
13:  $\sigma_{b}^{2}\leftarrow$ sample variance of $\{b_{i}\}$
14:end for
Implementation and Results
We implemented the GNMM in R using the gnmm.sgd and gnmm.predict routines provided in the Supplementary Material of [17] and tailored for our own dataset. Following the strategy of [17], the final visit for every subject was held out, so the training set comprised 4 9604\,960 records and the test set $915$. We compare the following models:
  • 1‑layer GNMM: one hidden layer with three ReLU nodes, ridge penalty $\lambda=0.001$, learning rate $0.005$, random intercept included.
  • 2‑layer GNMM: two hidden layers (three and two nodes), $\lambda=0.002$, learning rate $0.005$, random intercept included.
  • ANN baseline: one hidden layer with three nodes, $\lambda=0.001$, learning rate $0.001$, no random effect.
Evaluation.
Predictive accuracy was measured on the held‑out visits using mean squared error (MSE) and mean absolute error (MAE). Each model was trained 50times with independent random seeds; Table 6 reports the average test performance.
Table 4: Average test‑set error over five independent runs
Model MSE MAE
1‑layer GNMM $96.82$ $6.96$
2‑layer GNMM $106.09$ $7.56$
ANN (no random effect) $114.20$ $8.47$
The single‑layer GNMM attains the lowest prediction error, reducing test‑set MSE by $15\%$ relative to the two‑layer variant and by $15.3\%$ relative to the feed‑forward network without random effects.
3.5 Neural Mixed-Effects (NME) Model for Longitudinal UPDRS Prediction
Another recently introduced neural network model that can be applied to our case is the Neural Mixed-Effects (NME) model.
Classical mixed-effects models (LMM, GLMM) effectively handle subject heterogeneity in longitudinal data but are typically restricted to linear fixed effects. Conversely, standard neural networks can learn rich nonlinear relationships but often ignore the within-subject correlation inherent in repeated measures. The Neural Mixed Effects (NME) framework, as proposed by Wörtwein et al. [27], elegantly combines these strengths. This framework permits the inclusion of nonlinear subject-specific parameters at any layer of the network and utilizes stochastic gradient descent for optimization, which ensures scalability with both the number of patients ($m$) and the total number of visits (i.e., $\sum_{i=1}^{m}n_{i}$).
Applying the NME approach to our Parkinson’s tele-monitoring study offers several advantages. First of all, it allows for the learning of complex relationships between voice features and disease severity without the need for pre-specifying interaction terms.
Additionally, the NME model employs partial pooling for its parameter estimates. This approach allows the model to share information across different patients, leading to more robust and reliable individual-specific parameters, particularly for patients with fewer observations, by balancing individual data with overall population trends.
3.6 NME Parameterization.
Let $i=1,\dots,m$ index the $m=42$ participants in the Parkinson’s tele-monitoring study, and $j=1,\dots,n_{i}$ index their repeated visits. At each visit $j$ for participant $i$, we observe the response $Y_{ij}$, representing the UPDRS score, and a $p$-dimensional predictor vector $\mathbf{X}_{ij}\in\mathbb{R}^{17}$ which is consisted of test time (time of assessment) and 16 scaled voice features. Following Wörtwein et al. [27], the NME model decomposes the network parameters into two components:
  1. 1.
    A person-generic component $\bar{\bm{\theta}}$, which is shared across all participants and captures common trends.
  2. 2.
    A person-specific component $\bm{\eta}_{i}$, unique to participant $i$, capturing individual deviations from the generic trend.
The effective parameters for participant $i$ are thus $\bm{\theta}_{i}=\bar{\bm{\theta}}+\bm{\eta}_{i}$. The person-specific components $\bm{\eta}_{i}$ are typically regularized by assuming they follow a multivariate normal distribution, $\bm{\eta}_{i}\sim N(\mathbf{0},\bm{\Sigma})$, where $\bm{\Sigma}$ is a covariance matrix (often diagonal, e.g., $\bm{\Sigma}=\tau^{2}\mathbf{I}$).
Network Architecture for UPDRS Prediction.
For predicting total_UPDRS, we implement a two-hidden-layer multilayer perceptron (MLP) with $k_{1}=32$ units in the first hidden layer and $k_{2}=16$ units in the second hidden layer. The parameters of this network are decomposed into person-generic components (elements of $\bar{\bm{\theta}}$) and person-specific deviations (elements of $\bm{\eta}_{i}$) as defined in Equation (LABEL:eq:NME_param_split_combined). Specifically, the network operations involving these decomposed parameters are defined as:
$\bm{\alpha}^{(1)}_{ij}$ $=g_{1a}\!\left((\bar{\bm{\Omega}}^{(1)}+\bm{\eta}_{\Omega^{(1)},i})\mathbf{X}_{ij}+(\bar{\bm{\delta}}^{(1)}+\bm{\eta}_{\delta^{(1)},i})\right)$ (16)
$\bm{\alpha}^{(2)}_{ij}$ $=g_{1b}\!\left((\bar{\bm{\Omega}}^{(2)}+\bm{\eta}_{\Omega^{(2)},i})\bm{\alpha}^{(1)}_{ij}+(\bar{\bm{\delta}}^{(2)}+\bm{\eta}_{\delta^{(2)},i})\right)$ (17)
$\hat{Y}_{ij}\equiv\mu_{ij}^{\text{NME}}$ $=g_{0}\!\left((\bar{\bm{\omega}}^{(0)}+\bm{\eta}_{\omega^{(0)},i})\bm{\alpha}^{(2)}_{ij}+(\bar{\delta}^{(0)}+\bm{\eta}_{\delta^{(0)},i})\right)$ (18)
In these equations, $\mathbf{X}_{ij}$ represents the input features for subject $i$ at visit $j$. The term $\bm{\alpha}^{(1)}_{ij}$ denotes the activations of the first hidden layer. These are computed using the input-to-first-hidden-layer person-generic weights $\bar{\bm{\Omega}}^{(1)}$ (a component of $\bar{\bm{\theta}}$) and person-specific weight deviations $\bm{\eta}_{\Omega^{(1)},i}$ (a component of $\bm{\eta}_{i}$), as well as the corresponding biases $\bar{\bm{\delta}}^{(1)}$ and $\bm{\eta}_{\delta^{(1)},i}$, followed by the activation function $g_{1a}(\cdot)$. Similarly, $\bm{\alpha}^{(2)}_{ij}$ represents the activations of the second hidden layer, taking $\bm{\alpha}^{(1)}_{ij}$ as input. This layer uses person-generic weights $\bar{\bm{\Omega}}^{(2)}$ and biases $\bar{\bm{\delta}}^{(2)}$, with their respective person-specific deviations $\bm{\eta}_{\Omega^{(2)},i}$ and $\bm{\eta}_{\delta^{(2)},i}$, followed by its activation function $g_{1b}(\cdot)$. For both hidden layers, the activation function used is the Rectified Linear Unit (ReLU). The final prediction, $\hat{Y}_{ij}$ (or $\mu_{ij}^{\text{NME}}$), is obtained from the output layer. This layer takes $\bm{\alpha}^{(2)}_{ij}$ as input and applies the second-hidden-to-output-layer person-generic weights $\bar{\bm{\omega}}^{(0)}$ and output biases $\bar{\delta}^{(0)}$, along with their person-specific deviations $\bm{\eta}_{\omega^{(0)},i}$ and $\bm{\eta}_{\delta^{(0)},i}$. The output layer activation function $g_{0}(\cdot)$ is the identity function ($g_{0}(x)=x$), as total_UPDRS is a continuous response. Collectively, the person-generic parameters are $\bar{\bm{\theta}}=(\bar{\bm{\Omega}}^{(1)},\bar{\bm{\delta}}^{(1)},\bar{\bm{\Omega}}^{(2)},\bar{\bm{\delta}}^{(2)},\bar{\bm{\omega}}^{(0)},\bar{\delta}^{(0)})$, and the person-specific deviations for subject $i$ are $\bm{\eta}_{i}=(\bm{\eta}_{\Omega^{(1)},i},\bm{\eta}_{\delta^{(1)},i},\bm{\eta}_{\Omega^{(2)},i},\bm{\eta}_{\delta^{(2)},i},\bm{\eta}_{\omega^{(0)},i},\bm{\eta}_{\delta^{(0)},i})$. All these parameters are estimated during training. If a specific parameter (or an entire layer’s parameters) is not intended to have a patient-specific component, its corresponding entries in $\bm{\eta}_{i}$ are fixed at zero.
Loss Function and Optimization.
The NME objective function is optimized per epoch. For our regression task with squared error loss, $l(Y_{ij},\hat{Y}_{ij})=\frac{1}{2}(Y_{ij}-\hat{Y}_{ij})^{2}$, and assuming a diagonal person-specific parameter covariance $\bm{\Sigma}=\tau^{2}\mathbf{I}$ (implying $\bm{\Sigma}^{-1}=(1/\tau^{2})\mathbf{I}$), the objective function, adapted from Equation (1) of Wörtwein et al. [27], is:
$\mathcal{L}(\bar{\bm{\theta}},\{\bm{\eta}_{i}\})=\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\frac{1}{\sigma^{2}}\frac{1}{2}\left(Y_{ij}-\mu_{ij}^{\text{NME}}\right)^{2}+\sum_{i=1}^{m}\bm{\eta}_{i}^{\top}\bm{\Sigma}^{-1}\bm{\eta}_{i}$ (19)
where $\sigma^{2}$ is the observational (residual) variance, typically estimated as the mean squared error on the training data after each epoch. The first term encourages fidelity to the data, while the second term penalizes large deviations of person-specific parameters $\bm{\eta}_{i}$ from zero, effectively shrinking them towards the person-generic parameters $\bar{\bm{\theta}}$.
For stochastic gradient descent using mini-batches, the loss for a mini-batch $\mathcal{B}$ of size $B_{\text{batch}}$ (containing observations from a set of unique subjects $\mathcal{B}_{\text{subjects}}$) is formulated. The data fidelity part is the average loss over the batch. The regularization penalty is applied per subject within the batch, scaled by the proportion of that subject’s total observations present in the current batch, as described by Wörtwein et al. [27]. Thus, the mini-batch loss is:
$\tilde{\mathcal{L}}_{\mathcal{B}}=\frac{1}{B_{\text{batch}}}\sum_{(i,j)\in\mathcal{B}}\left[\frac{1}{\sigma^{2}}\frac{1}{2}\left(Y_{ij}-\mu_{ij}^{\text{NME}}\right)^{2}\right]+\sum_{k\in\mathcal{B}_{\text{subjects}}}\left(\frac{N_{k,\mathcal{B}}}{m_{k}}\bm{\eta}_{k}^{\top}\bm{\Sigma}^{-1}\bm{\eta}_{k}\right)$ (20)
where $N_{k,\mathcal{B}}$ is the number of observations for subject $k$ in the current mini-batch $\mathcal{B}$, and $m_{k}$ is the total number of training observations for subject $k$. This scaling ensures that the regularization for each subject is weighted according to its representation in the batch relative to its total contribution.
Gradient Updates.
The parameters $(\bar{\bm{\theta}},\{\bm{\eta}_{i}\}_{i=1}^{m})$ are updated using gradients derived from the loss function $\mathcal{L}$ (Equation (19)). For any parameter $\phi$ (which could be a component of $\bar{\bm{\theta}}$ or a component of some $\bm{\eta}_{k}$), the update uses its partial derivative.
Let $E_{ij}=Y_{ij}-\mu_{ij}^{\text{NME}}$ be the prediction error for subject $i$ at visit $j$. The derivative of the data fidelity part of the loss with respect to the model output $\mu_{ij}^{\text{NME}}$ (assuming squared error loss $l(Y,\hat{Y})=\frac{1}{2}(Y-\hat{Y})^{2}$) is $\frac{\partial l}{\partial\mu_{ij}^{\text{NME}}}=-(Y_{ij}-\mu_{ij}^{\text{NME}})=-E_{ij}$. Thus, the common error signal propagated back from the loss, scaled by the residual variance, is:
$\delta_{ij}^{\text{out}}=-\frac{E_{ij}}{\sigma^{2}}$
For our defined two-hidden-layer network, where $\mu_{ij}^{\text{NME}}$ is defined by equations (16), (17), and (18), we have $g_{0}^{\prime}(x)=1$ (identity output activation), and $g_{1a}^{\prime}(\cdot)$ and $g_{1b}^{\prime}(\cdot)$ are the derivatives of the ReLU activation functions for the first and second hidden layers, respectively. Let $\bm{s}^{(1)}_{ij}$ and $\bm{s}^{(2)}_{ij}$ be the pre-activations for the first and second hidden layers.
The gradients for the weight parameters are derived as follows. These equations illustrate how the error signal is backpropagated and combined with local inputs/activations to update each parameter.
Output Layer Parameters:
The output layer directly computes the prediction $\mu_{ij}^{\text{NME}}$. For an element $p$ of the generic output weights $\bar{\bm{\omega}}^{(0)}$ (connecting $p$-th unit of the second hidden layer to the output):
$\frac{\partial\mathcal{L}}{\partial\bar{\omega}^{(0)}_{p}}=\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\delta_{ij}^{\text{out}}\cdot\alpha^{(2)}_{ij,p}$
This gradient term aggregates the product of the output error signal and the corresponding activation from the second hidden layer across all observations. For the $p$-th element of a person-specific deviation of an output weight $\bm{\eta}_{\omega^{(0)},k}$ for subject $k$:
$\frac{\partial\mathcal{L}}{\partial\eta_{\omega^{(0)}_{p},k}}=\sum_{j=1}^{n_{k}}\delta_{kj}^{\text{out}}\cdot\alpha^{(2)}_{kj,p}+2[\bm{\Sigma}^{-1}\bm{\eta}_{k}]_{\omega^{(0)}_{p}}$
where the second term serve as a regularization term that penalizes large deviations.
Second Hidden Layer Parameters:
Gradients for the second hidden layer involve backpropagating the error signal through the output layer weights. For an element $\bar{\Omega}^{(2)}_{pc}$ of the generic weights of the second hidden layer (connecting $c$-th unit of the first hidden layer to $p$-th unit of the second hidden layer):
$\frac{\partial\mathcal{L}}{\partial\bar{\Omega}^{(2)}_{pc}}=\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\delta_{ij}^{\text{out}}\cdot(\bar{\omega}^{(0)}_{p}+\eta_{\omega^{(0)}_{p},i})\cdot g_{1b}^{\prime}(s^{(2)}_{ij,p})\cdot\alpha^{(1)}_{ij,c}$
Here, the error signal is weighted by the effective output weight and the derivative of the second hidden layer’s activation, then multiplied by the activation from the first hidden layer. For an element $\eta_{\Omega^{(2)}_{pc},k}$ of a person-specific deviation of a second hidden layer weight for subject $k$:
$\frac{\partial\mathcal{L}}{\partial\eta_{\Omega^{(2)}_{pc},k}}=\sum_{j=1}^{n_{k}}\delta_{kj}^{\text{out}}\cdot(\bar{\omega}^{(0)}_{p}+\eta_{\omega^{(0)}_{p},k})\cdot g_{1b}^{\prime}(s^{(2)}_{kj,p})\cdot\alpha^{(1)}_{kj,c}+2[\bm{\Sigma}^{-1}\bm{\eta}_{k}]_{\Omega^{(2)}_{pc}}$
First Hidden Layer Parameters:
Gradients for the first hidden layer involve further backpropagation through the second hidden layer weights. For an element $\bar{\Omega}^{(1)}_{cl}$ of the generic weights of the first hidden layer (connecting $l$-th input feature to $c$-th unit of the first hidden layer):
$\frac{\partial\mathcal{L}}{\partial\bar{\Omega}^{(1)}_{cl}}=\sum_{i=1}^{m}\sum_{j=1}^{n_{i}}\left(\sum_{p=1}^{k_{2}}\delta_{ij}^{\text{out}}\cdot(\bar{\omega}^{(0)}_{p}+\eta_{\omega^{(0)}_{p},i})\cdot g_{1b}^{\prime}(s^{(2)}_{ij,p})\cdot(\bar{\Omega}^{(2)}_{pc}+\eta_{\Omega^{(2)}_{pc},i})\right)g_{1a}^{\prime}(s^{(1)}_{ij,c})\cdot X_{ij,l}$
where $k_{2}$ is the number of units in the second hidden layer.
For an element $\eta_{\Omega^{(1)}_{cl},k}$ of a person-specific deviation of a first hidden layer weight for subject $k$:
$\frac{\partial\mathcal{L}}{\partial\eta_{\Omega^{(1)}_{cl},k}}=$ $\sum_{j=1}^{n_{k}}\left(\sum_{p=1}^{k_{2}}\delta_{kj}^{\text{out}}\cdot(\bar{\omega}^{(0)}_{p}+\eta_{\omega^{(0)}_{p},k})\cdot g_{1b}^{\prime}(s^{(2)}_{kj,p})\cdot(\bar{\Omega}^{(2)}_{pc}+\eta_{\Omega^{(2)}_{pc},k})\right)g_{1a}^{\prime}(s^{(1)}_{kj,c})\cdot X_{kj,l}$
$+2[\bm{\Sigma}^{-1}\bm{\eta}_{k}]_{\Omega^{(1)}_{cl}}$
The pre-activations are $s^{(1)}_{ij,c}$ for the $c$-th unit of the first hidden layer, and $s^{(2)}_{ij,p}$ for the $p$-th unit of the second hidden layer for observation $(i,j)$.
Gradients for all bias terms ($\bar{\bm{\delta}}^{(1)},\bm{\eta}_{\delta^{(1)},i},\bar{\bm{\delta}}^{(2)},\bm{\eta}_{\delta^{(2)},i},\bar{\delta}^{(0)},\bm{\eta}_{\delta^{(0)},i}$) follow analogously by applying the chain rule, where the input to the bias is 1. During mini-batch optimization, these sums are taken over the observations $(i,j)$ in the current mini-batch $\mathcal{B}$, and the regularization term’s gradient is applied only for subjects $k$ whose parameters $\bm{\eta}_{k}$ are being updated.
Summary of the Algorithm
The Neural Mixed Effects (NME) model is trained using an iterative, optimization-based procedure, which involves employing stochastic gradient descent (e.g., Adam optimizer) within each epoch to update the person-generic parameters $\bar{\bm{\theta}}$ and the person-specific deviations $\{\bm{\eta}_{i}\}_{i=1}^{m}$. The variance components, namely the observational (residual) variance $\sigma^{2}$ and the covariance matrix of the person-specific parameters $\bm{\Sigma}$, are generally updated between epochs. For instance, $\sigma^{2}$ can be estimated based on the mean squared error from the training data using the current parameter estimates. The covariance matrix $\bm{\Sigma}$ is often assumed to be diagonal (e.g., $\bm{\Sigma}=\tau^{2}\mathbf{I}$) for scalability and is updated based on the sample covariance of the current person-specific deviations.
This iterative training procedure allows the NME model to learn both the overall population trend (via $\bar{\bm{\theta}}$) and subject-specific nonlinear deviations (via $\bm{\eta}_{i}$) simultaneously. The person-specific deviations are regularized by their prior distribution, typically governed by the estimated covariance structure $\bm{\Sigma}$, which helps prevent overfitting and allows for robust estimation even for subjects with limited data.
The overall iterative process is outlined below:
Algorithm 2 General Training Procedure for the Neural Mixed Effects (NME) Model
1:Scaled training features $\mathbf{X}^{\prime}=\{\mathbf{X}^{\prime}_{ij}\}$, Scaled training targets $Y^{\prime}=\{Y^{\prime}_{ij}\}$, Subject indices for observations.
2:Architectural choices (e.g., number of layers, units), Number of epochs $E$, Learning rate $\eta$, Batch size $B$.
3:Initialize person-generic parameters $\bar{\bm{\theta}}$ (e.g., Xavier initialization).
4:Initialize person-specific deviations $\{\bm{\eta}_{i}\}_{i=1}^{m}$ (e.g., to zeros or small random values).
5:Initialize covariance matrix $\bm{\Sigma}$ (e.g., as a scaled identity matrix).
6:Initialize residual variance $\sigma^{2}$ (e.g., to 1 or based on an initial pass over the data).
7:Initialize optimizer (e.g., Adam with learning rate $\eta$).
8:for epoch = 1 to $E$ do
9:  Shuffle training data $(\mathbf{X}^{\prime},Y^{\prime})$.
10:  for each batch $b$ of size $B$ do
11:   For each observation $(k,j)$ in batch $b$ (subject $k$, observation $j$):
12:    Compute prediction $\hat{Y}^{\prime}_{kj}=f(\mathbf{X}^{\prime}_{kj};\bar{\bm{\theta}}+\bm{\eta}_{k})$.
13:   Compute mini-batch loss $\tilde{\mathcal{L}}_{b}$ (e.g., based on Eq. (20)), using current $\sigma^{2},\bm{\Sigma}$.
14:   Compute gradients w.r.t. $\bar{\bm{\theta}}$ and relevant $\{\bm{\eta}_{k}\}$ for subjects in the batch: $\nabla_{\bar{\bm{\theta}}}\tilde{\mathcal{L}}_{b}$, $\nabla_{\bm{\eta}_{k}}\tilde{\mathcal{L}}_{b}$.
15:   Update $\bar{\bm{\theta}}$ and relevant $\{\bm{\eta}_{k}\}$ using the optimizer step.
16:  end for
17:  Update $\sigma^{2}$ based on the average squared residuals over the full training set using current $\bar{\bm{\theta}},\{\bm{\eta}_{i}\}$.
18:  Update $\bm{\Sigma}$ based on the sample covariance of the current person-specific deviations $\{\bm{\eta}_{i}\}_{i=1}^{m}$.
19:  Adjust learning rate or check for early stopping criteria if applicable.
20:end for
21:return Learned parameters $\hat{\bar{\bm{\theta}}}$, $\{\hat{\bm{\eta}}_{i}\}$, $\hat{\bm{\Sigma}}$, $\hat{\sigma}^{2}$.
Implementation and Results
Our application of the Neural Mixed Effects (NME) model to predict Total UPDRS scores was based on the publicly available PyTorch implementation provided by Wörtwein et al. [27] and tailored for our case.
Input voice features and test_time were standardized. We configured the NME model with a two-hidden-layer MLP (32 units in the first layer, 16 in the second, both using ReLU activations) as the base network, applying person-specific random effects ($\bm{\eta}_{i}$) to the output layer’s bias. The model was trained for 4000 epochs using the Adam optimizer and a batch size of 512. During training, the observational variance $\sigma^{2}$ and a diagonal person-specific parameter covariance $\bm{\Sigma}=\tau^{2}\mathbf{I}$ were updated iteratively, consistent with the NME framework.
Predictive accuracy was measured using mean squared error (MSE) and mean absolute error (MAE). Table 5 summarizes the key performance metrics.
Table 5: Performance of the NME Model on the Test Set for Total UPDRS Prediction.
Model MSE MAE
NME-MLP 103.4075 8.1786
4 Analysis of the Results
Table 6: Predictive performance (MSE and MAE) of six models. Test set consists of each subject’s last test time point (42 total), with the remaining observations used for training.
Model MSE MAE
LMM $7.70$ $2.25$
GAMM $\mathbf{6.56}$ $\mathbf{2.00}$
1‑layer GNMM $96.82$ $6.96$
2‑layer GNMM $106.09$ $7.56$
NME-MLP $103.41$ $8.18$
ANN (no random effect) $114.20$ $8.47$
In this study, we developed and compared a suite of modeling approaches for predicting Parkinson’s disease progression, using a rich longitudinal voice dataset and the total UPDRS score as the clinical outcome. We evaluated both traditional statistical models, Linear Mixed Effects Models (LMM) and Generalized Additive Mixed Models (GAMM), as well as machine learning-based extensions, Generalized Neural Network Mixed Models (GNMM) and Neural Mixed Effect Models (NME-MLP).
To assess predictive performance, we constructed a test set composed of each subject’s last available time point (42 in total), while the remaining data were used for model training. This setup reflects a realistic clinical use case: forecasting future UPDRS values for already-observed patients, rather than for entirely new individuals.
Table 6 reports the mean squared error (MSE) and mean absolute error (MAE) of each model. Among all methods, GAMM achieved the best performance with the lowest MSE ($6.56$) and MAE ($2.00$), indicating that the spline-based temporal effect captured meaningful nonlinear disease progression patterns. The LMM, although simpler, performed nearly as well (MSE = $7.70$), confirming the value of mixed-effects modeling with carefully selected covariates and interactions.
In contrast, the deep learning models, 1-layer and 2-layer GNMMs, ANN without random effects, and NME-MLP—performed substantially worse, with MSEs exceeding $96$ and MAEs exceeding 6, which is counter-intuitive. This finding is consistent with our recent evidence that larger, more complex biomedical deep models do not automatically improve clinical prediction over simpler baselines when the dataset and signal structure do not support the added capacity [29]. As we normally assume that newer and complicated models outperforms the order and simpler ones. But that is not always the case, for any datasets which have many observations but only a modest number of predictors ($n\gg p$) a simple linear or spline-based model can already approximate the input–output mapping well, so the added capacity of deep networks does not translate into lower error unless it is strongly regularized. While these architectures are expressive, their complexity and lack of explicit structure for within-subject correlation hinder their predictive accuracy in settings like ours.
5 Summary and Future Work
In summary, traditional statistical models, especially GAMM, outperformed more complex neural models for the near-term prediction of the severity of Parkinson’s disease. Our findings indicate that incorporating smooth effects and subject-level random structures remains a robust and interpretable strategy, particularly when sample size is limited and the goal is individualised disease monitoring.
Another key limitation of the current neural approaches is their lack of an explicit variable-selection mechanism. Neither the GNMM nor the NME-MLP papers explain how to decide which acoustic features should enter the model or how to discard variables that do not carry predictive information. In our experiments we chose the input set by trial and error, which is slow and can leave many irrelevant predictors in the network. Retaining uninformative variables inflates estimation variance, hides the contribution of important features, and complicates clinical interpretation. Building on the benchmark guidelines summarised by Tong et al. [22], we plan to construct a transparent test bed that compares classical statistical models with neural mixed-effects alternatives.
Future research should therefore focus on automatic variable selection within neural mixed-effects models. One option is to place sparsity-inducing penalties, such as the $\ell_{1}$ or group lasso, on the first-layer weights so that inputs with small effects are driven towards zero and effectively removed. Bayesian spike-and-slab priors offer another alternative that yields posterior inclusion probabilities for each variable. Recent knockoff-based screens, adapted to longitudinal data, could provide finite-sample false-discovery control while allowing the network to remain flexible. Another route is to pre-screen variables with traditional parametric methods, then fine-tune a compact neural model that employs only the retained subset. These strategies would give clinicians a clearer picture of which voice markers matter, while reducing overfitting and computation time.
Another important concern is whether the proposed models can be embedded in a telemedicine workflow for continuous, home-based monitoring. A practical system would capture speech with low-cost microphones or smartphone sensors, stream the audio to a secure server, and run the GAMM or neural mixed-effects model in near real time. When bandwidth is limited, edge devices could perform feature extraction locally and transmit only the compressed acoustic markers, reducing latency and preserving patient privacy. Model predictions would then feed into a clinician dashboard that issues alerts if the estimated Total UPDRS exceeds a patient-specific threshold or shows a sustained upward trend. Achieving this vision requires (i) robust noise-handling and calibration procedures so that recordings from different hardware remain comparable, (ii) lightweight model implementations that can update with incremental data without full retraining, and (iii) clear data-governance protocols that comply with HIPAA and related regulations. Pilot studies integrating these components will be essential to demonstrate feasibility, user acceptance, and clinical utility before large-scale deployment.More broadly, the same push toward clinically usable, end-to-end AI systems connects telemonitoring with other healthcare pipelines such as drug discovery and delivery, where data quality, interpretability, and deployment constraints also shape real-world impact [30].
SUPPORTING INFORMATION
Web Appendices referenced in Sections 3 and 4 are provided in the Supporting Information. Python and R code, along with a simulated example, are available at https://github.com/RanTongUTD/Parkinson-Prediction/.